Circular No. 8890

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 17P/HOLMES C. Salyk and G. A. Blake, California Institute of Technology; M. J.

Mumma, B. P. Bonev, M. A. DiSanti, and G. L. Villanueva, Goddard Space Flight Center; Y. Radeva, University of Maryland; K. Magee-Sauer, Rowan University; and E. L. Gibb, University of Missouri, St. Louis, report that they observed comet 17P/Holmes on Oct. 29 and 30 UT (r = 2.46 AU) using NIRSPEC at the W. M. Keck Observatory. On Oct. 29, H_{2}O (eleven lines, 2.9-μm hot bands), C_{2}H_{6} (eight Q-branches, 3.35-μm ν_{7} band), C_{2}H_{2} (four lines, 3.0-μm ν_{3} band), HCN (nine lines, 3.0-μm ν_{3} band), and CH_{3}OH (3.52-μm P-, Q-, R-branches, ν_{3} band) were detected. Rotational temperatures for H_{2}O, HCN, and C_{2}H_{6} were ∼ 60 K. All lines were strongly peaked on the nucleus region, as expected for parent volatiles. Preliminary production rates (× 10^{27} molecules/s) are: H_{2}O, 275; C_{2}H_{6}, 5.9; C_{2}H_{2}, 1.5; HCN, 1.5; and CH_{3}OH, 11. The abundance ratios are: H_{2}O : C_{2}H_{6} : C_{2}H_{2} : HCN : CH_{3}OH = 100 : 2.2 : 0.54 : 0.54 : 4.0. Comet 17P seems to be enriched in organics, like C/2001 A2. On Oct. 30, abundance ratios for H_{2}O, C_{2}H_{2}, and HCN were similar to those on Oct. 29. The R0 and P2 lines of CO (4.7-μm 1-0 band) were detected; if T_{rot} = 60 K, the formal CO abundance is 14 (± 4) relative to water (100).

COMETS C/2007 M5–M9 (SOHO) Further to IAUC 8889, additional near-sun comets have been found

on SOHO website images — C/2007 M5 and C/2007 M8 being non-group, and the rest being Kreutz sungrazers. C/2007 M5, which has the smallest perihelion distance on record (q = 0.0011 AU), entered the C2 field-of-view at peak brightness (mag ∼ 6.5) and faded rapidly, disappearing from view before it reached the occulting disk. C/2007 M6 was stellar in appearance (mag ∼ 7) in C3 images, but diffuse with no tail in C2 images. C/2007 M7 and C/2007 M9 were faint (mag ∼ 8) and diffuse. C/2007 M8 was very small, circular (nearly stellar in appearance), and faint (mag ∼ 7.5–8), appearing to be marginally diffuse.

Comet 2007 UT α_{2000} δ_{2000} Inst. F MPEC
C/2007 M5 June 25.229 6^{h}06.^{m}0 +22^{o}37′ C2  BZ 2007-U15
C/2007 M6  25.346 6 05.4 +21 38 C3/2 BZ 2007-U15
C/2007 M7  25.413 6 07.9 +21 39 C2  BZ 2007-U16
C/2007 M8  25.564 6 17.8 +22 32 C2  HS 2007-U16
C/2007 M9  27.604 6 16.5 +21 44 C2  BZ 2007-U16
2007 November 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green